\begin{thebibliography}{10}

\bibitem{AHI2025}
{\sc Annala, T., Hoyois, M., and Iwasa, R.}
\newblock Algebraic cobordism and a {Conner}-{Floyd} isomorphism for algebraic
  {K}-theory.
\newblock {\em J. Am. Math. Soc. 38}, 1 (2025), 243--289.

\bibitem{AMMN}
{\sc Antieau, B., Mathew, A., Morrow, M., and Nikolaus, T.}
\newblock On the {Beilinson} fiber square.
\newblock {\em Duke Math. J. 171}, 18 (2022), 3707--3806.

\bibitem{bhatt2022absolute}
{\sc Bhatt, B., and Lurie, J.}
\newblock Absolute prismatic cohomology.
\newblock {\em arXiv:2201.06120\/} (2022).

\bibitem{BhattMathew2023}
{\sc Bhatt, B., and Mathew, A.}
\newblock Syntomic complexes and {{\(p\)}}-adic {\'e}tale {Tate} twists.
\newblock {\em Forum Math. Pi 11\/} (2023), 26.
\newblock Id/No e1.

\bibitem{BhattMorrowScholze2019}
{\sc Bhatt, B., Morrow, M., and Scholze, P.}
\newblock Topological {H}ochschild homology and integral {$p$}-adic {H}odge
  theory.
\newblock {\em Publ. Math. Inst. Hautes \'{E}tudes Sci. 129\/} (2019),
  199--310.

\bibitem{BEK14}
{\sc Bloch, S., Esnault, H., and Kerz, M.}
\newblock {$p$}-adic deformation of algebraic cycle classes.
\newblock {\em Invent. Math. 195}, 3 (2014), 673--722.

\bibitem{BlochKato1986}
{\sc Bloch, S., and Kato, K.}
\newblock {$p$}-adic \'etale cohomology.
\newblock {\em Inst. Hautes \'Etudes Sci. Publ. Math.}, 63 (1986), 107--152.

\bibitem{BO74}
{\sc Bloch, S., and Ogus, A.}
\newblock Gersten's conjecture and the homology of schemes.
\newblock {\em Ann. Sci. \'Ecole Norm. Sup. (4) 7\/} (1974), 181--201 (1975).

\bibitem{CHK97}
{\sc Colliot-Th\'el\`ene, J.-L., Hoobler, R.~T., and Kahn, B.}
\newblock The {B}loch-{O}gus-{G}abber theorem.
\newblock In {\em Algebraic {$K$}-theory ({T}oronto, {ON}, 1996)}, vol.~16 of
  {\em Fields Inst. Commun.} Amer. Math. Soc., Providence, RI, 1997,
  pp.~31--94.

\bibitem{CSS83}
{\sc Colliot-Th\'el\`ene, J.-L., Sansuc, J.-J., and Soul\'e, C.}
\newblock Torsion dans le groupe de {C}how de codimension deux.
\newblock {\em Duke Math. J. 50}, 3 (1983), 763--801.

\bibitem{Elkik1973}
{\sc Elkik, R.}
\newblock Solutions d'\'{e}quations \`a coefficients dans un anneau
  hens\'{e}lien.
\newblock {\em Ann. Sci. \'{E}cole Norm. Sup. (4) 6\/} (1973), 553--603 (1974).

\bibitem{ElmantoMorrow}
{\sc Elmanto, E., and Morrow, M.}
\newblock Motivic cohomology of equicharacteristic schemes.
\newblock {\em arXiv:2309.08463\/} (2023).

\bibitem{Fu02}
{\sc Fujiwara, K.}
\newblock A proof of the absolute purity conjecture (after {G}abber).
\newblock In {\em Algebraic geometry 2000, {A}zumino ({H}otaka)}, vol.~36 of
  {\em Adv. Stud. Pure Math.} Math. Soc. Japan, Tokyo, 2002, pp.~153--183.

\bibitem{GabberAffineAnalogue}
{\sc Gabber, O.}
\newblock {{\(K\)}}-theory of {Henselian} local rings and {Henselian} pairs.
\newblock In {\em Algebraic \(K\)-theory, commutative algebra, and algebraic
  geometry. Proceedings of the US-Italy joint seminar, held June 18-24, 1989 at
  Santa Margherita Ligure, Italy, with support from the National Science
  Foundation and Consiglio Nazionale delle Ricerche}. Providence, RI: American
  Mathematical Society, 1992, pp.~59--70.

\bibitem{Gabber1994}
{\sc Gabber, O.}
\newblock Gersten's conjecture for some complexes of vanishing cycles.
\newblock {\em Manuscr. Math. 85}, 3-4 (1994), 323--343.

\bibitem{Geisser2004}
{\sc Geisser, T.}
\newblock Motivic cohomology over {D}edekind rings.
\newblock {\em Math. Z. 248}, 4 (2004), 773--794.

\bibitem{GeisserLevine2000}
{\sc Geisser, T., and Levine, M.}
\newblock The {$K$}-theory of fields in characteristic {$p$}.
\newblock {\em Invent. Math. 139}, 3 (2000), 459--493.

\bibitem{GiLe87}
{\sc Gillet, H., and Levine, M.}
\newblock The relative form of {G}ersten's conjecture over a discrete valuation
  ring: the smooth case.
\newblock {\em J. Pure Appl. Algebra 46}, 1 (1987), 59--71.

\bibitem{GrosSuwa1988}
{\sc Gros, M., and Suwa, N.}
\newblock La conjecture de {Gersten} pour les faisceaux de {Hodge}-{Witt}
  logarithmique. ({The} {Gersten} conjecture for the logarithmic {Hodge}-{Witt}
  sheaves).
\newblock {\em Duke Math. J. 57}, 2 (1988), 615--628.

\bibitem{SGA4}
{\sc Grothendieck, A., and Verdier, J.-L.}
\newblock {\em Th\'eorie des topos et cohomologie \'etale des sch\'emas. {T}ome
  1: {T}h\'eorie des topos}.
\newblock Lecture Notes in Mathematics, Vol. 269. Springer-Verlag, Berlin-New
  York, 1972.
\newblock S\'eminaire de G\'eom\'etrie Alg\'ebrique du Bois-Marie 1963--1964
  (SGA 4), Dirig\'e par M. Artin, A. Grothendieck, et J. L. Verdier. Avec la
  collaboration de N. Bourbaki, P. Deligne et B. Saint-Donat.

\bibitem{Ha66}
{\sc Hartshorne, R.}
\newblock {\em Residues and duality}.
\newblock Lecture notes of a seminar on the work of A. Grothendieck, given at
  Harvard 1963/64. With an appendix by P. Deligne. Lecture Notes in
  Mathematics, No. 20. Springer-Verlag, Berlin-New York, 1966.

\bibitem{JS}
{\sc Jannsen, U., and Saito, S.}
\newblock Kato conjecture and motivic cohomoogy over finite fields.
\newblock {\em Preprint\/}.

\bibitem{KaS86}
{\sc Kato, K., and Saito, S.}
\newblock Global class field theory of arithmetic schemes.
\newblock In {\em Applications of algebraic {$K$}-theory to algebraic geometry
  and number theory, {P}art {I}, {II} ({B}oulder, {C}olo., 1983)}, vol.~55 of
  {\em Contemp. Math.} Amer. Math. Soc., Providence, RI, 1986, pp.~255--331.

\bibitem{Kerz2010}
{\sc Kerz, M.}
\newblock Milnor {$K$}-theory of local rings with finite residue fields.
\newblock {\em J. Algebraic Geom. 19}, 1 (2010), 173--191.

\bibitem{KerzEsnaultWittenberg2016}
{\sc Kerz, M., Esnault, H., and Wittenberg, O.}
\newblock A restriction isomorphism for cycles of relative dimension zero.
\newblock {\em Camb. J. Math. 4}, 2 (2016), 163--196.

\bibitem{Lu16}
{\sc L{\"u}ders, M.}
\newblock On a base change conjecture for higher zero-cycles.
\newblock {\em Homology Homotopy Appl. 20}, 1 (2018), 59--68.

\bibitem{Lueders2019}
{\sc L{\"u}ders, M.}
\newblock Algebraization for zero-cycles and the $p$-adic cycle class map.
\newblock {\em Mathematical Research Letters 26}, 2 (2019), 557--585.

\bibitem{Lu2021}
{\sc L{\"u}ders, M.}
\newblock Bloch-{Ogus} theory for smooth and semi-stable schemes in mixed
  characteristic.
\newblock {\em arXiv:2201.11553\/} (2022).

\bibitem{LuMo20}
{\sc L{\"u}ders, M., and Morrow, M.}
\newblock Milnor {{\(K\)}}-theory of {{\(p\)}}-adic rings.
\newblock {\em J. Reine Angew. Math. 796\/} (2023), 69--116.

\bibitem{Milne1986}
{\sc Milne, J.~S.}
\newblock Values of zeta functions of varieties over finite fields.
\newblock {\em Am. J. Math. 108\/} (1986), 297--360.

\bibitem{Quillen1973}
{\sc Quillen, D.}
\newblock Higher algebraic {$K$}-theory. {I}.
\newblock In {\em Algebraic {$K$}-theory, {I}: {H}igher {$K$}-theories ({P}roc.
  {C}onf., {B}attelle {M}emorial {I}nst., {S}eattle, {W}ash., 1972)\/} (1973),
  Lecture Notes in Math., Vol. 341, Springer, Berlin, pp.~85--147.

\bibitem{SS14}
{\sc Saito, S., and Sato, K.}
\newblock Zero-cycles on varieties over {$p$}-adic fields and {B}rauer groups.
\newblock {\em Ann. Sci. \'Ec. Norm. Sup\'er. (4) 47}, 3 (2014), 505--537.

\bibitem{sakagaito2023etale}
{\sc Sakagaito, M.}
\newblock On \'{e}tale hypercohomology of henselian regular local rings with
  values in $p$-adic \'{e}tale tate twists.
\newblock {\em arXiv:2002.04797\/} (2023).

\bibitem{Sa07}
{\sc Sato, K.}
\newblock {$p$}-adic \'etale {T}ate twists and arithmetic duality.
\newblock {\em Ann. Sci. \'Ecole Norm. Sup. (4) 40}, 4 (2007), 519--588.
\newblock With an appendix by Kei Hagihara.

\bibitem{Sa13}
{\sc Sato, K.}
\newblock Cycle classes for {$p$}-adic \'etale {T}ate twists and the image of
  {$p$}-adic regulators.
\newblock {\em Doc. Math. 18\/} (2013), 177--247.

\bibitem{SS18}
{\sc Schmidt, J., and Strunk, F.}
\newblock Stable {$\Bbb A^1$}-connectivity over {D}edekind schemes.
\newblock {\em Ann. K-Theory 3}, 2 (2018), 331--367.

\bibitem{SS19}
{\sc Schmidt, J., and Strunk, F.}
\newblock A {Bloch}-{Ogus} theorem for henselian local rings in mixed
  characteristic.
\newblock {\em Math. Z. 303}, 4 (2023), 24.
\newblock Id/No 82.

\bibitem{Schneider1994}
{\sc Schneider, P.}
\newblock {{\(p\)}}-adic points of motives.
\newblock In {\em Motives. Proceedings of the summer research conference on
  motives, held at the University of Washington, Seattle, WA, USA, July
  20-August 2, 1991}. Providence, RI: American Mathematical Society, 1994,
  pp.~225--249.

\end{thebibliography}

\documentclass[a4paper, 10pt, twoside]{article}

\usepackage{amsmath, amscd, amsfonts, amssymb, amsthm, latexsym, url, color, todonotes, bm, framed, rotating, enumerate} %pdflscape}
\usepackage{graphicx}
\usepackage[left=1in, right=1in, top=1.2in, bottom=1in, includefoot, headheight=13.6pt]{geometry}
\usepackage{mathtools}
\input{xy}
\xyoption{all}

\usepackage[colorlinks,breaklinks=true]{hyperref}
\usepackage[figure,table]{hypcap}
\hypersetup{
	bookmarksnumbered,
	pdfstartview={FitH},
	citecolor={black},%{blue},
	linkcolor={black},%{red},
	urlcolor={black},
	pdfpagemode={UseOutlines}
}
\makeatletter
\newcommand\org@hypertarget{}
\let\org@hypertarget\hypertarget
\renewcommand\hypertarget[2]{%
  \Hy@raisedlink{\org@hypertarget{#1}{}}#2%
} 
\makeatother 

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{Key lemma}[theorem]{Key lemma}

\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}

\newtheorem{definition}[theorem]{Definition}
\newtheorem{remark}[theorem]{Remark}

\newtheorem{conjecture}[theorem]{Conjecture}

\renewcommand{\labelenumi}{(\roman{enumi})}

\newcommand{\xysquare}[8]{
\[\xymatrix{
#1 \ar@{#5}[r] \ar@{#6}[d] & #2 \ar@{#7}[d]\\
#3 \ar@{#8}[r] & #4
}\]
}

\newcommand{\bb}{\mathbb}

\newcommand{\comment}[1]{}

\renewcommand{\phi}{\varphi}

\newcommand{\roi}{\mathcal{O}}

\newcommand{\ul}[1]{\underline{#1}}

\newcommand{\xto}{\xrightarrow}

\renewcommand{\cal}{\mathcal}
\renewcommand{\hat}{\widehat}

\renewcommand{\tilde}{\widetilde}

\renewcommand{\projlim}{\varprojlim}

\DeclareMathOperator{\Frac}{Frac}

\DeclareMathOperator{\Spec}{Spec}

\newcommand{\CH}{C\!H}

\def\T{{\cal T}}
\def\et{{\mathrm{\acute{e}t}}} 
\newcommand{\K}{\hat{\cal K}}

\usepackage[bbgreekl]{mathbbol}
\DeclareSymbolFontAlphabet{\mathbbm}{bbold}

\usepackage{fancyhdr}

\usepackage{sectsty}
\sectionfont{\Large\sc\centering}
\chapterfont{\large\sc\centering}
\chaptertitlefont{\LARGE\centering}
\partfont{\centering}

\begin{document}
\itemsep0pt

\title{The Gersten conjecture for $p$-adic \'etale Tate twists and the $p$-adic cycle class map}

\author{Morten L\"uders}
%\address{Institut de Mathématiques de Jussieu–Paris Rive Gauche, UMPC - 4 place Jussieu, Case 247, 75252 Paris, France}
%\email{morten.lueders@ur.de}
%\classno{}

\date{}
%\extraline{}

\maketitle

\begin{abstract}
We prove the Gersten conjecture for $p$-adic \'etale Tate twists for a smooth scheme $X$ in mixed characteristic in the Nisnevich topology. Our main observation is that, while $p$-adic \'etale Tate twists are not $\bb A^1$-invariant, for the proof  of the Gersten conjecture it suffices that they satisfy the $\bb P^1$-bundle formula. This fits nicely with the emphasis on the projective bundle formula in non $\bb A^1$-invariant motivic cohomology recently developed by Elmanto-Morrow and Annala-Hoyois-Iwasa. Furthermore, identifying $p$-adic \'etale Tate twists with the syntomic cohomology defined by Bhatt-Morrow-Scholze, the result generalises the Gersten conjecture for logarithmic deRham-Witt sheaves due to Gros-Suwa to arbitrary characteristic. In the second part of the article, we revisit the cycle class map from thickened zero-cycles on the special fiber of $X$ to \'etale cohomology with coefficients in $p$-adic \'etale Tate twists previously studied in \cite{Lueders2019}. This cycle class map is important in the study of zero-cycles on smooth projective varieties over local fields and the approach to the cycle class map which we use in this article is more conceptual and, in contrast to the approach in \textit{loc. cit.}, works for arbitrary finite residue fields.
\end{abstract}

%\tableofcontents

\section{Introduction}
Let $k$ be a field and $X$ a smooth $k$-scheme of dimension $d$. Then Quillen \cite{Quillen1973} showed that the Gersten complex
$$0\rightarrow \mathcal{K}_{q,X}\xrightarrow{} \bigoplus_{x\in X^{(0)}}i_{x*} K_q(x)\rightarrow \bigoplus_{x\in X^{(1)}} i_{x*}K_{q-1}(x)\rightarrow...\to \bigoplus_{x\in X^{(d)}}i_{x*}K_{q-d}(x)\to 0 $$
for algebraic $K$-theory is exact in the Zariski topology, thereby verifying Gersten's conjecture for algebraic $K$-theory for smooth schemes over a field. This has had, amoong others, many applications to algebraic cycles because of the Bloch-Quillen formula $H^q(X,\mathcal{K}_{q,X})\cong \CH^q(X)$ and has motivated the study of the Gersten conjecture for other theories. For example, the Gersten conjecture may be formulated for any cohomology theory with supports $H^{*}(-): \mathrm{Sch}_S\to \rm Ab$ and base scheme $S$.\footnote{A cohomology theory with supports is a contravariant functor, which to any triple $Z\subset Y\subset X$ with $Z,Y$ closed in $X$ associates a long exact sequence
$$\dots \to H^*_Z(X,\cal T_r(n))\to H^*_Y(X,\cal T_r(n))\to H^*_{Y-Z}(X-Z,\cal T_r(n))\to H^{*+1}_Z(X,\cal T_r(n))\to \dots$$
These sequences give rise to the coniveau spectral sequence, whose rows are the Cousin complexes (see \cite[Sec. 5]{CHK97} or Section \ref{section_general_reductions}).}
\begin{conjecture}(Gersten) Let $\mathcal{H}^{q}_{\mathrm{Zar},X}$ be the Zariski sheafification of the presheaf $U\mapsto H^q(U)$. If $X$ is a regular scheme, then the Cousin complex
$$\bigoplus_{x\in X^{(0)}}i_{x*} H^{q}_x(X)\rightarrow \bigoplus_{x\in X^{(1)}} i_{x*}H^{q+1}_x(X)\rightarrow...\to \bigoplus_{x\in X^{(d)}}i_{x*}H^{q+d}_x(X)\to 0 $$
is a resolution of $\mathcal{H}^{q}_{\mathrm{Zar},X}$ in the Zariski topology.
\end{conjecture}
Examples over a field include $H^{q}(-)=H^{q}(-,\mu_\ell^n)$ with $(\ell,ch(k))=1$ and $H^{q}(-)=H^{q}(-,W_r\Omega^n_{X,\log})$ in which case $\mathcal{H}^{q}_{\mathrm{Zar},X}$ is isomorphic to $R^q\epsilon_* \mu_\ell^n$ and $R^q\epsilon_*W_r\Omega^n_{\log}$ respectively. Here $\epsilon:X_{\et}\to X_{\rm Zar}$ is the natural morphism of sites. In these cases, the cohomology of the respective Cousin complexes are the $E_2$-terms of the Leray spectral sequence associated to $\epsilon$. The main ingredient of Quillen's proof is a noether normalisation lemma to reduce the question to $\bb A^d_k$ and this idea is also used by Bloch-Ogus \cite{BO74} to show that the  Gersten conjecture holds for smooth schemes over a field for \'etale cohomology with coefficients prime to the characteristic of $k$, i.e. $H^{q}(-)=H^{q}(-,\mu_\ell^n)$ with $(\ell,ch(k))=1$. In the approach of Bloch-Ogus, however, difficult properties like Poincar\'e duality are required. Subsequently, Gabber developed a geometric presentation lemma and observed that this lemma and the $\bb P^1$-bundle formula suffice to prove the Gersten conjecture. This was most notably used by Gros-Suwa \cite{GrosSuwa1988} to show that the Gersten conjecture holds for $H^{q}(-)=H^{q}(-,W_n\Omega^n_{\log})$. The logarithmic deRham-Witt sheaves $W_n\Omega^n_{\log}$ are particularly important since they can be used to define $p$-adic \'etale motivic cohomology in characteristic $p$. Gabber's approach was axiomatised by Colliot-Th\'el\`ene-Hoobler-Kahn \cite{CHK97}. 
More recently, in the context of their motivic cohomology of equicharacteristic schemes, Elmanto-Morrow have shown that any presheaf on smooth $k$-schemes which satisfies the $\bb P^1$-bundle formula is deflatable and that any such Nisnevich presheaf which is deflatable satisfies the Gersten injectivity \cite[Lem. 6.11]{ElmantoMorrow} (see also Remark \ref{remark_Elmanto_Morrow}).

In this article, we focus on the mixed characteristic situation: let $A$ be a discrete valuation ring with local parameter $\pi$, perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero. Let $S=\Spec(A)$ and $X$ be a smooth $S$-scheme of dimension $d$. In this situation, the Gersten conjecture for algebraic K-theory with finite coefficients has been proved by Gillet-Levine \cite{GiLe87} and Geisser-Levine \cite{GeisserLevine2000} and for motivic cohomology with finite coefficients by Geisser \cite{Geisser2004}. The case of \'etale cohomology with torsion coefficients prime to $ch(k)$ has been studied by Schmidt-Strunk \cite{SS19} and the author \cite{Lu2021}. All of these cases have in common, that the theories are $\bb A^1$-invariant, at least for regular schemes. 
Our focus in this article will be on the Gersten conjecture for $p$-adic \'etale Tate twists which we denote by $\cal T_r(n)\in D^b(X_{\et},\bb Z/p^r)$. $p$-adic \'etale Tate twists were first defined by Schneider in the smooth case and then by Sato in the semi-stable case \cite{Sa07} and are generally considered to be the right $p$-adic \'etale motivic cohomology with mod $p^r$-coefficients in mixed characteristic for smooth and semi-stable schemes.
% and thereby closely related to $p$-adic algebraic K-theory. 
The first main result of this article is the following:
\begin{theorem}\label{theorem_intro}(Thm. \ref{theorem_main_in_text})
Assume that $A$ is henselian and let $\epsilon:X_{\et}\to X_{\rm Nis}$ be the natural morphism of sites. Then for all $n\geq 0$ the complex of sheaves
$$0\to R^q\epsilon_*\cal T_r(n) \to \bigoplus_{x\in X^{(0)}}i_{*,x}H^{q}_x(X,\cal T_r(n))\to \bigoplus_{x\in X^{(1)}}i_{*,x}H^{q+1}_x(X,\cal T_r(n))\to $$
$$\dots\to \bigoplus_{x\in X^{(d)}}i_{*,x}H^{q+d}_x(X,\cal T_r(n))\to 0 $$
is exact in the Nisnevich topology.
\end{theorem}
The cohomology of $p$-adic \'etale Tate twists, just like the cohomology of logarithmic deRham-Witt sheaves, is not $\bb A^1$-invariant but satisfies the projective bundle formula (see Section \ref{section_proj_bun_formula}) and our key observation, going back to Gabber's mentioned above, is that the $\bb P^1$-bundle formula in combination with a geometric presentation lemma due to Schmidt-Strunk suffices to prove Theorem \ref{theorem_intro}. The theorem may therefore be considered to be an analogue of the main theorem of Gros-Suwa on the Gersten conjecture for logarithmic deRham-Witt sheaves.
Furthermore, both theories have found a common generalisation in the syntomic complexes $\bb Z/p^r(n)_X$ first defined by Bhatt-Morrow-Scholze \cite{BhattMorrowScholze2019} for quasi-syntomic rings and whose definition has been extended to arbitrary schemes by Bhatt-Lurie \cite{bhatt2022absolute}: for any regular $\bb F_p$-scheme $X$ there exist isomorphisms $\bb Z/p^r(n)_X\cong W_n\Omega^n_{X,\log}[-n]$ \cite[Sec. 8]{BhattMorrowScholze2019} and $\bb Z/p^r(n)_X\cong \T_r(n)\in D^b(X_{\et},\bb Z/p^r)$ for $X$ as in Theorem \ref{theorem_intro} \cite[Thm. 5.8] {BhattMathew2023}. In Theorem \ref{theorem_intro}, the reader may therefore replace $\T_r(n)$ by $\bb Z/p^r(n)_X$ and $A$ by a perfect field or a henselian dvr with perfect residue field. Due to the importance of syntomic cohomology, we focus on this case, for an axiomatisation of the approach, however, see Remark \ref{remark_Elmanto_Morrow}. Finally, we remark that these observations fit nicely with the emphasis on the projective bundle formula in non $\bb A^1$-invariant motivic cohomology recently developed by Elmanto-Morrow \cite{ElmantoMorrow} and Annala-Hoyois-Iwasa \cite{AHI2025}.

%This theorem complements other results on the Gersten conjecture for smooth schemes in mixed characteristic. Gillet-Levine prove it for $K$-theory \cite{GiLe87}, Geisser for motivic cohomology \cite{Geisser2004}, Schmidt-Strunk \cite{SS19} and the author \cite{Lu2021} for \'etale cohomology with finite coefficients away from the residue characteristic. For $q\leq n$ our theorem can also be deduced from Geisser's results for motivic cohomology and his results on the Beilinson-Lichtenbaum conjecture in mixed characteristic. The approach of this article just uses properties of $p$-adic \'etale Tate twists. The key observation is the following. One cannot use the formalism of a cohomology theory with duality combined with Quillen's trick as in \cite{BO74,Lu2021} since there is no Poincar\'e duality in all weights for $p$-adic \'etale Tate twists. Therefore, one has to use Gabber's approach \cite{Gabber1994,CHK97}, the main ingredients of which are a geometric presentation lemma, generalised to mixed characteristic by Schmidt-Strunk \cite{SS18}, and strict $\bb A^1$-invariance or the projective bundle formula. While $p$-adic \'etale Tate twists are not strictly $\bb A^1$-invariant, they satisfy the projective bundle formula.

Our second main theorem is the following:
\begin{theorem}(Thm. \ref{theorem_comparison})\label{theorem_2}
Let the notation be as above but let $X/S$ be smooth projective and $d$ be the relative dimension of $X$ over $S$. Assume that $k$ is finite. Let $X_k$ be the special fiber of $X$ (in the following also denoted by $X_1$) and denote the inclusion by $i:X_k\to X$. Let $X_n=X\times_A A/\pi^n$ and let $\hat{\cal K}^M$ denote the improved Milnor K-sheaf defined in \cite{Kerz2010}. 
Then for $q\in\{0,1\}$ there are isomorphisms
$$H^{d-q}_{\rm Nis}(X_k,i^*\hat{\cal K}^M_{d,X}/p^r) \xto{\cong}  \projlim_n H^{d-q}_{\rm Nis}(X_k,\hat{\cal K}^M_{d,X_n}/p^r)\xto{\cong} H^{2d-q}_\et(X_k,i^*\cal T_r(d)). $$
\end{theorem}
This was proved for $q=0$ and $A=W(k)$ and assuming that the $k$ is sufficiently large in \cite[Prop. 1.4]{Lueders2019} and the proof used the difficult techniques of \cite[Sec. 7]{BEK14} which only work if $ch(k)>d$. The proof we give in this article is more conceptual and works for arbitrary finite residue fields. The key ingredients are the calculation of the top cohomology of $\cal T_r(n)$, which also enters in the proof of Theorem \ref{theorem_intro}, the Kato conjectures for logarithmic deRham-Witt sheaves and the Gersten conjecture for $\hat{\cal K}^M_{d,X,\rm Nis}/p^r$ which was recently proved by Morrow and the author.
The importance of Theorem \ref{theorem_2} comes from the commutative diagram  
$$
\xymatrix{
 \CH^d(X_K)/p^r & \CH^d(X)/p^r\ar[r]^-{res^d_{/p^r}} \ar[d]_{cl_X} \ar@{->>}[l] & \projlim_n H^{d}_{\rm Nis}(X_k,\K^M_{d,X_n}/p^r) \ar[d]^{cl_{X_k}}  \\
& H^{2d}_\et(X, \cal T_r(d)) \ar[r]^-\cong & H^{2d}_\et(X_k,i^*\cal T_r(d)),   \\
}
$$
$K=\Frac(A)$, which is used in \cite{Lueders2019} and \cite{SS14} to study the Chow group of zero cycles mod $p^r$ of smooth projective schemes over local fields. Furthermore, $cl_{X_k}$ can be considered to be a thickened analogue of the isomorphism $\CH^d(X_k)/p^r\xto{\cong} H^{2d}(X,W_r\Omega^d_{r,\log}[-d])\cong \pi_1^{\rm ab}(X_k)/p^r$ which is studied in higher-dimensional unramified class field theory (mod $p^r$) over finite fields. 

\paragraph{Acknowledgement.} I would like to thank Tess Bouis for helpful comments and acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 \textit{Geometry and Arithmetic of Uniformized Structures}, project number 444845124. Furthermore, I would like to thank an anonymous referee for a very detailed and helpful report.

\section{$p$-adic \'etale Tate twists}\label{section_etaletatetwists}
Let $A$ be a discrete valuation ring with perfect residue field of characteristic $p>0$ and fraction field $K$ of characteristic zero. Let $X$ be a regular scheme which is flat of finite type over $S=\Spec A$ and which is a smooth or semistable family over $S$. Let $Y$ be the special fiber and $X_K$ the generic fiber with inclusions
$$X_K\xto{j} X \xleftarrow{i} Y.$$
In \cite{Sa07}, and for all $n\geq 0$ and $r\geq 1$, Sato constructs a pair $(\T_r(n)_X,t')$, with $\T_r(n)_X\in D^b(X_{\et},\bb Z/p^r\bb Z)$ unique up to unique isomorphism, fitting into a distinguished triangle of the form
$$i_*\nu^{n-1}_{Y,r}[-n-1]\xto{} \T_r(n)_X\xto{t'}\tau_{\leq n}Rj_*\mu_{p^r}^{\otimes n}\xto{}i_*\nu^{n-1}_{Y,r}[-n].$$
The complexes $\T_r(n)_X$ are called \textit{$p$-adic \'etale Tate twists}.
If $X$ is smooth over $S$, then $\nu^{n-1}_{Y,r}=W_r\Omega^{n-1}_{Y,\log}$ and $\nu^{<0}_{Y,r}=0$. In the smooth case the objects $\T_r(n)_X$ were first defined by Schneider \cite{Schneider1994} and extensively studied by Geisser \cite{Geisser2004}.
Following \cite{Sa07}, we list the properties of $p$-adic \'etale Tate twists which we need for the Gersten conjecture.

\begin{enumerate}
\item  
There is an isomorphism
       $t : j^*\T_r(n)_X \simeq \mu_{p^r}^{\otimes n}$ (see \cite[Def. 4.2.4]{Sa07}).
\item %{\bf T2 (Acyclicity).}
$\T_r(n)_X$ is concentrated in $[0,n]$, i.e.,
   the $q$-th cohomology sheaf is zero unless $0 \leq q \leq n$ (see \cite[Def. 4.2.4]{Sa07}).
\item %{\bf T3 (Purity).}
For a locally closed regular subscheme $i : Z \hookrightarrow X$ of codimension
          $c \,(\ge 1)$ which is contained in the special fiber $Y$, there is a Gysin isomorphism
$W_r\Omega_{Z,\log}^{n-c}[-n-c] \xto{\cong}\tau_{\leq n+c} R i^!\T_r(n)_X$ in $D^b(Z_{\et},\bb Z/p^r\bb Z)$ (see \cite[Thm. 4.4.7]{Sa07}).
\item %{\bf T5 (Product structure).}
There is a unique morphism
$
\T_r(m)_X \otimes^{\bb L} \T_r(n)_X
      \to \T_r(m+n)_X$ in $D^b(X_{\et},\bb Z/p^r\bb Z)$
that extends the natural isomorphism
         $\mu_{p^r}^{\otimes m} \otimes \mu_{p^r}^{\otimes n}
           \simeq \mu_{p^r}^{\otimes m+n}$ on $X[1/p]$ (see \cite[Prop. 4.2.6]{Sa07}).
\item Let $X'$ be another scheme which is flat of finite type over $S$
    and for which the objects $\T_r(n)_{X'}$ are defined.
Let $f : X' \to X$ be a morphism of schemes and let
    $\psi : X'[1/p] \to X[1/p]$ be the induced morphism. Then there is a unique morphism
$$
\begin{CD}
f^*\T_r(n)_X
       @>>> \T_r(n)_{X'} 
\end{CD}
$$
in $D^b(X'_{\et},\bb Z/p^r\bb Z)$ that extends the natural isomorphism
    $\psi^*\mu_{p^r}^{\otimes n} \simeq \mu_{p^r}^{\otimes n}$
     on $X'[1/p]$ (see \cite[Prop. 4.2.8]{Sa07}). We denote the induced map on cohomology by $$f^*:H^q(X,\T_r(n)_X)\to H^q(X',f^*\T_r(n)_X)\to H^q(X',\T_r(n)_{X'}).$$

Assume now that $f$ is proper, and put
    $c:=\dim(X)-\dim(X')$. Then by \cite[Thm. 7.1.1]{Sa07} there is a unique morphism
$$
\begin{CD}
   Rf_*\T_r(n-c)_{X'}[-2c]
       @>>> \T_r(n)_X
\end{CD}
$$
in $D^b(X_{\et},\bb Z/p^r\bb Z)$ that extends the trace morphism
   $R\psi_*\mu_{p^r}^{\otimes n-c}[-2c] \to \mu_{p^r}^{\otimes n}$
     on $X[1/p]$. The latter trace morphism (on the generic fiber) was first defined by Deligne in \cite[XVIII, Thm. 2.9]{SGA4} under some additional assumptions. Using Gabber's absolute purity theorem \cite{Fu02}, Deligne's trace can be extended to proper morphisms.
   \end{enumerate}  
%\end{quote}
%\smallskip
%\noindent

\begin{proposition}\label{proposition_purity_tate_twists} Let $X$ be as above and $i : Z \hookrightarrow X$ a regular closed subscheme of codimension
          $c \,(\ge 1)$ which is contained in the special fiber $Y$. Then there are isomorphisms
$$H^{n+c}_Z(X,\T_r(n))\cong H^{n-c}(Z,W_r\Omega^{n-c}_{Z,\log}[-(n-c)])$$
for all $n\geq 0$ (by definition, $W_r\Omega^{q}_{Z,\log}$ is the zero sheaf for $q<0$).
%In particular, for $x\in X$ a point of characteristic $p$,
%$$H^{c+n}_x(X,\T_r(n))\cong H^{n-c}(x,W_r\Omega^{n-c}_{x,\log}[-(n-c)])=:  H^{n-c}(x,\bb Z/p^r(n-c)).$$
%\todo{replace $W_r$ by $v$.}
\end{proposition}
\begin{proof}
Consider the Grothendieck spectral sequence
$$E_2^{u,v}=H^{u}(Z,R^vi^!\T_r(n))\Longrightarrow R^{v+u}(\Gamma_Z\circ i^!)\T_r(n)=:H^{v+u}_Z(X,\T_r(n)).$$
By purity (property (iii) above), this spectral sequence looks as follows:
$$\xymatrix{
n+c+1  & H^{0}(Z,R^1i^!\T_r(n)) & \dots & & \\
n+c  & H^0(Z,W_r\Omega^{n-c}_{Z,\log}) \ar[drr] & H^1(Z,W_r\Omega^{n-c}_{Z,\log}) & \dots  &\\
n+c-1  & 0 & 0 & 0 & \dots
}$$
This implies that $H^0(Z,W_r\Omega^{n-c}_{Z,\log})\cong H^{n+c}_Z(X,\T_r(n))$. 

%For $x\in X$ as in the proposition, passing to the colimit over all neighbourhoods $U$ of $x$, we get that
%$$H^{c+n}_x(X,\T_r(n)):=\indlim_{U} H^{c+n}_{\overline{\{x\}}\cap U}(U,\T_r(n))\cong $$
%$$\indlim_{V\subset \overline{\{x\}}}H^{n-c}(V,W_r\Omega^{n-c}_{V,\log}[-(n-c)])=:H^{n-c}(x,W_r\Omega^{n-c}_{x,\log}[-(n-c)]).$$
\end{proof}

\subsection{The projective bundle formula}\label{section_proj_bun_formula}
For the projective bundle formula we need to extend the definition of $p$-adic \'etale Tate twists to negative weights.
\begin{definition}(\cite[Def. 3.5(i)]{Sa13})
Let the notation be as above. For $n<0$ and $r\geq 1$, we define $\T_r(n):=j_!\mathcal{H}om_{X_K} (\mu_{p^r}^{\otimes -n},\bb Z/p^r)[0]$ placed in degree zero in $D^b(X_{\et},\bb Z/p^r)$.
\end{definition}
Let $X$ be as in the beginning of the section and $E$ be a vector bundle of rank $a+1$ on $X$ and $f:\bb P(E)\to X$ be the associated projective bundle. Let $\roi(1)_E$ be the tautological invertible sheaf on $\bb P(E)$. By \cite[4.5.1]{Sa07} there is a distinguished triangle
$$\cal T_r(1)_{\bb P(E)}\to \roi^\times_{\bb P(E)}[0]\xto{\times p^r} \roi^\times_{\bb P(E)}[0]\to \cal T_r(1)_{\bb P(E)}[1]$$
in $D^b(\bb P(E)_{\et})$ and we denote the image of $\roi(1)_E$ under the induced map
$$H^1(\bb P(E),\roi^\times)\to H^2(\bb P(E),\cal T_r(1)_{\bb P(E)})$$
by $\xi$.
The composition
$$\cal T_r(n-q)_{X}[-2q]\xto{f^*} Rf_*\cal T_r(n-q)_{\bb P(E)}[-2q]\xto{-\cup \xi^q}Rf_*\cal T_r(n)_{\bb P(E)},$$
$0\leq q \leq a$, where $f^*$ is the adjoint of the morphism defined in (v), induces a canonical morphism
$$\gamma_E:\bigoplus_{q=0}^a \cal T_r(n-q)_{X}[-2q]\to Rf_*\cal T_r(n)_{\bb P(E)}$$
in $D^b(X_{\et},\bb Z/p^r)$.
\begin{proposition}\label{proposition_projective_bundle_formula} \cite[Thm. 4.1]{Sa13} \quad $\gamma_E$ is an isomorphism for all $n\in \bb Z$.
\end{proposition}
For a more general statement concerning syntomic cohomology see also \cite[Thm. 9.1.1]{bhatt2022absolute}. As mentioned in the introduction, the syntomic complexes of \textit{loc. cit.} coincide with Sato's $p$-adic \'etale Tate twists for schemes for which the latter are defined by \cite[Thm. 5.8]{BhattMathew2023}.

\section{General reductions for the Gersten conjecture}\label{section_general_reductions}
Let $A$ be a discrete valuation ring with local parameter $\pi$, perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero and $S=\Spec(A)$. 
Let $\cal S_S$ be the category of regular semistable schemes which are flat of finite type over $S=\Spec A$. Let  $\cal P_S$ be the category of pairs $(X,Z)$, where $X\in \cal S_S$ and $Z$ is a closed subset of $X$. A morphism $f:(X',Z')\to (X,Z)$ in $\cal P_S$ consists of a morphism $f:X'\to X$ in $\cal S_S$ such that $f^{-1}(Z)\subset Z'$.
We define a cohomology theory with supports by the contravariant functor
$$(X,Z)\mapsto H^*_Z(X,\cal T_r(n))$$
from $\cal P_S$ to the category of abelian groups.
For any triple $Z\subset Y\subset X$ with $Z,Y$ closed in $X$ there is a long exact sequence
$$\dots \to H^*_Z(X,\cal T_r(n))\to H^*_Y(X,\cal T_r(n))\to H^*_{Y-Z}(X-Z,\cal T_r(n))\to H^{*+1}_Z(X,\cal T_r(n))\to \dots$$
(\cite[Sec. 1.9]{Sa07}).
Furthermore, we have excision: let $Z\subset X$ and $Z'\subset X'$ be closed subschemes and $\pi:X'\to X$ be an \'etale morphism such that the restriction of $\pi$ to $Z'$ is an isomorphism and $\pi(X'-Z')\subset X-Z$, then 
$$H^*_Z(X,\cal T_r(n))\xto{\cong}H^*_{Z'}(X',\pi^*\cal T_r(n)).$$
In the following let $\tau\in \{\rm Zar,\rm Nis\}$ denote the Zariski or Nisnevich topology. Let $\cal H^*(\cal T_r(n))$ be the sheafification of the presheaf $U\mapsto H^*(U,\cal T_r(n))$ for the $\tau$-topology on $X$.
\begin{definition}
\renewcommand{\labelenumi}{(\arabic{enumi})}
\begin{enumerate}
\item Let $X\in \cal S_S$. We let
$$Z^p(X):=\{Z\subset X\; |\;Z \;\mathrm{ closed},\; \mathrm{codim}_X(Z)\geq p\}$$
for all $p\in \bb N$ and let $Z^p(X)=\emptyset$ for $q<0$. If $X$ is clear from the context, then we also write $Z^p$ instead of $Z^p(X)$. Let $Z^p/Z^{p+1}$ denote the ordered set of pairs $(Z, Z')\in Z^p\times Z^{p+1}$ such that $Z'\subset Z$, with the ordering $(Z,Z')\geq (Z_1,Z_1')$ if $Z\supset Z_1$ and $Z'\supset Z_1'$. The group $Z^p(X)$ should not be confused with the group of algebraic cycle, i.e. the free abelian group generated by integral closed subschemes of codimension $p$ of $X$, which is the standard notation. Since we do not use the latter group elsewhere in the article, we stick to the notation used by Bloch-Ogus in \cite[Sec. 3]{BO74}.
\item For the $\tau$-topology on $X$, we define $\cal H^q_{Z^p}(\cal T_r(n))$ to be the sheaf associated to the presheaf
$$U\mapsto \varinjlim_{Z\in Z^p(U)} H^q_{Z}(U,\cal T_r(n))$$
and $\cal H^q_{Z^p/Z^{p+1}}(\cal T_r(n))$ to be the sheaf associated to the presheaf
$$U\mapsto \varinjlim_{(Z,Z')\in Z^p(U)/Z^{p+1}(U)} H^q_{Z-Z'}(U-Z',\cal T_r(n)).$$
\item Let
%$$H_q(x,n):=\varinjlim_{U\subseteq \overline{\{x\}}} H_q(U,n),$$
%$$H^q(x,\cal T_r(n)):=\varinjlim_{U\subseteq \overline{\{x\}}} H^q(U,\cal T_r(n))$$
%and
$$H^q_x(X,\cal T_r(n)):=\varinjlim_{U\ni x} H^q_{\overline{\{x\}}\cap U}(U,\cal T_r(n)),$$
where the colimit is taken over all open subschemes 
%$U$ in $\overline{\{x\}}$ and all open subschemes 
$U$ in $X$ containing $x$.
\end{enumerate}
\end{definition}
\begin{proposition}\label{proposition_cousin_complex}
Let $\epsilon:X_{\et}\to X_{\tau}$ be the natural morphism of sites. If the natural map $$(*)\colon\cal H^q_{Z^{p+1}}(\cal T_r(n))\to \cal H^q_{Z^p}(\cal T_r(n))$$ is zero (in the $\tau$-topology) for all $p,q,n\in \mathbb{N}$, then there exists an exact sequence of $\tau$-sheaves   
$$
0\to R^q\epsilon_*\cal T_r(n) \to \bigoplus_{x\in X^{(0)}}i_{*,x}H^{q}_x(X,\cal T_r(n))\to \bigoplus_{x\in X^{(1)}}i_{*,x}H^{q+1}_x(X,\cal T_r(n))\to $$
 $$\dots\to \bigoplus_{x\in X^{(d)}}i_{*,x}H^{q+d}_x(X,\cal T_r(n))\to 0.  
$$ 
\end{proposition}
\begin{proof}
Assuming $(*)=0$ for all $p,q,n\in \mathbb{N}$, the exact sequences 
$$\dots\to \cal H^{q-1}_{Z^p/Z^{p+1}}(\cal T_r(n))\to\cal H^q_{Z^{p+1}}(\cal T_r(n))\xto{0} \cal H^q_{Z^p}(\cal T_r(n))\to \cal H^{q}_{Z^p/Z^{p+1}}(\cal T_r(n))\to \dots$$
and
$$\dots\to \cal H^{q-1}_{Z^{p+1}/Z^{p+2}}(\cal T_r(n))\to\cal H^q_{Z^{p+2}}(\cal T_r(n))\xto{0} \cal H^q_{Z^{p+1}}(\cal T_r(n))\to \cal H^{q}_{Z^{p+1}/Z^{p+2}}(\cal T_r(n))\to \dots $$
splice together as follows:
$$\dots\to \cal H^{q-1}_{Z^p/Z^{p+1}}(\cal T_r(n))\twoheadrightarrow\cal H^q_{Z^{p+1}}(\cal T_r(n))\hookrightarrow \cal H^q_{Z^{p+1}/Z^{p+2}}(\cal T_r(n)) \twoheadrightarrow \cal H^q_{Z^{p+2}}(\cal T_r(n))\to\dots\; .$$
Note that the first two non-zero terms in the complex of the proposition coincide with $\cal H^q_{Z^{0}}(\cal T_r(n))$ and $\cal H^q_{Z^{0}/Z^{1}}(\cal T_r(n))$ and that $\cal H^{q-1}_{Z^{-1}/Z^{0}}(\cal T_r(n))=0$. It therefore remains to show that therefore there are isomorphisms 
$$\cal H^{q}_{Z^p/Z^{p+1}}(\cal T_r(n))\cong \bigoplus_{x\in X^{(q)}}i_{*,x}H^{q}_x(X,\cal T_r(n)).$$ 
 This follows from Zariski excision: if $T_1,...,T_r$ are pairwise disjoint closed subsets of $X$, then $$\bigoplus_i H^q_{T_i}(X,\cal T_r(n))\cong H^q_{\cup T_i}(X,\cal T_r(n))$$ (see \cite[Lemma  1.2.1]{CHK97}). 
\end{proof}

\begin{remark} The complexes starting with $\bigoplus_{x\in X^{(0)}}$ in Proposition \ref{proposition_cousin_complex} are called \textit{Cousin complexes} and coincide by the construction in the proof with the rows of the coniveau spectral sequence (see \cite[Sec. 1.1]{CHK97} or \cite[Ch. IV]{Ha66}).
\end{remark} 

The following definition is used in applications to show that $(*)$ is zero.
\begin{definition}
Let $X\in \cal S_S$ and $x\in X$ a point. We say that $\cal T_r(n)$ is effaceable at $x$ if the following condition is satisfied: Given $p\geq 0$, for any affine $\tau$-neighbourhood $W$ of $x$ and any closed subscheme $Z\in Z^{p+1}(W)$ there exists a $\tau$-neighbourhood $f:U\to W$ of $x$ and a closed subscheme $Z'\in Z^p(U)$ containing $f^{-1}(Z)$ and $x$ such that the composition 
$$H^q_{Z}(W,\cal T_r(n))\to H^q_{f^{-1}(Z)}(U,\cal T_r(n))\to H^q_{Z'}(U,\cal T_r(n))$$
is zero for all $q$.
\end{definition}

\section{Proof of the Gersten conjecture}
We fix the following notation. Let $A$ be a discrete valuation ring with local parameter $\pi$, perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero, and $S=\Spec(A)$. 
%Let $A$ be a discrete valuation ring with perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero. Let $S=\Spec(A)$. 

In order to prove our main theorem, the Gersten conjecture for $p$-adic \'etale Tate twists, we closely follow Gabber's strategy \cite{Gabber1994} as exposed in \cite{CHK97}. A key ingredient is the following lemma. In equal characteristic, it has appeared in \cite[Key lemma, Sec. 4.1]{CHK97}, \cite[p. 621]{GrosSuwa1988} and, more recently, in \cite[Lem. 6.12]{ElmantoMorrow} (see also Remark \ref{remark_Elmanto_Morrow}). 
\begin{lemma}\label{prop_commutativity}
Let $V$ be a smooth $S$-scheme, $p:\bb A^1_V\to V$ and $\tilde p:\bb P^1_V\to V$ the natural projections. Let $j: \bb A^1_V\to \bb P^1_V$ be the inclusion and $s_\infty:V\to \bb P^1_V$ the section of $\tilde p$ at infinity. Let $F$ be a closed subset of $V$. Then the diagram
$$\xymatrix{
 H^{q}_{\bb A^1_F}(\bb A^1_V,\cal T_r(n))  &  \\
 & H^{q}_F(V ,\cal T_r(n)) \ar[ul]_{p^*} & \\
  H^{q}_{\bb P^1_F}(\bb P^1_V,\cal T_r(n)) \ar[uu]^{j^*}  \ar[ur]^{s_\infty^*} & 
}$$
commutes.
\end{lemma}
\begin{proof}
By Proposition \ref{proposition_projective_bundle_formula} there is an isomorphism $H^{q}_{\bb P^1_F}(\bb P^1_V,\cal T_r(n))\cong H^{q-2}_F(V ,\cal T_r(n-1))\oplus  H^{q}_F(V ,\cal T_r(n))$. Since the restrictions $j^* \roi(1)$ and $s_\infty^* \roi(1)$ are trivial, this implies that $j^*|_{H^{q-2}_F(V ,\cal T_r(n-1))}=0$ and $s_\infty^*|_{H^{q-2}_F(V ,\cal T_r(n-1))}=0$ and therefore that the diagram commutes.
\end{proof}

\begin{proposition}\label{proposition_A1_effacement}
Let $V$ be a smooth $S$-scheme, $F$ a closed subset of $V$ and $F'$ a closed subset of $\bb A^1_F$ such that the projection $f:F'\to F$ is finite. Diagrammatically:
$$\xymatrix{
 F' \ar@{^{(}->}[r] \ar[rd]_f   & \bb A^1_F \ar[d] \ar@{^{(}->}[r]  & \bb A^1_V \ar[d]^p  \\
 & F \ar@{^{(}->}[r]  & V. 
}$$
Then the natural map 
$$ H^{q}_{F'}(\bb A^1_V,\cal T_r(n))\to H^{q}_{\bb A^1_F}(\bb A^1_V,\cal T_r(n)) $$
is zero.
\end{proposition}
\begin{proof} The argument is essentially the  same as in the proof of \cite[Thm. 4.2.1]{CHK97}.
Let $k:\bb P^1_V-F'\hookrightarrow \bb P^1_V$ be the open immersion. Since $s_\infty(V)\cap F'=\emptyset$ as $f$ is finite, $s_\infty$ can be factorised into $s_\infty=k\circ s'$.
The statement now follows from the following commutative diagram in which the isomorphism on the left is due to excision, the middle triangle commutes by Lemma \ref{prop_commutativity} and the lower horizontal composition is part of the localisation exact sequence and therefore zero.

$$\xymatrix{
 H^{q}_{F'}(\bb A^1_V,\cal T_r(n)) \ar[r]   & H^{q}_{\bb A^1_F}(\bb A^1_V,\cal T_r(n))  & & & \\
 &  \ar[u] & H^{q}_F(V ,\cal T_r(n)) \ar[ul]_{p^*} & & \\
H^{q}_{F'}(\bb P^1_V,\cal T_r(n))\ar[r] \ar[uu]_\cong  & H^{q}_{\bb P^1_F}(\bb P^1_V,\cal T_r(n)) \ar[uu]^{j^*} \ar[rr]^-{k^*} \ar[ur]^{s_\infty^*} & & H^{q}_{\bb P^1_F-F'}(\bb P^1_V-F' ,\cal T_r(n)). \ar[ul]_{s'^*} 
}$$
\end{proof}

\begin{theorem}\label{theorem_main_in_text}
Assume that $S$ is henselian. Let $X/S$ be a smooth scheme of dimension $d$ and let $\epsilon:X_{\et}\to X_{\rm Nis}$ be the natural morphism of sites.
The complex of sheaves
$$0\to R^q\epsilon_*\cal T_r(n) \to \bigoplus_{x\in X^{(0)}}i_{*,x}H^{q}_x(X,\cal T_r(n))\to \bigoplus_{x\in X^{(1)}}i_{*,x}H^{q+1}_x(X,\cal T_r(n))\to $$
$$\dots\to \bigoplus_{x\in X^{(d)}}i_{*,x}H^{q+d}_x(X,\cal T_r(n))\to 0 $$
is exact in the Nisnevich topology for all $n\geq 0$.
\end{theorem}
\begin{proof}
For the following argument with torsion coefficients of order prime to $p$ see \cite[Cor. 3.5]{Lu2021} and \cite[Sec. 5,6]{SS19}. We need to show that the natural map $$(*)\colon\cal H^q_{Z^{p+1}}(\cal T_r(n))\to \cal H^q_{Z^p}(\cal T_r(n))$$ is zero in the Nisnevich topology for all $p,q,n\in \mathbb{N}$. We show $(*)=0$ on stalks. Let $x\in X$. We first assume that $k$ is infinite.
Let $W$ be an affine Nisnevich neighbourhood of $x$ and $Z\in Z^{p+1}(W)$ and assume that $Z_k\neq W_k$, $Z_k$ and $W_k$ being the special fibers of $Z$ and $W$ respectively. Then by \cite[Thm. 2.4]{SS18} we can find a Nisnevich neighbourhood $f:U\to W$ of $x$ and an \'etale morphism
$$\rho=(\psi,\nu): U\to V\times\bb A^1_S\subset \bb A^{d-1}_S\times \bb A^1_S,$$
where $V$ is an open subset of $\bb A^{d-1}_S$, with the following properties: Let $Z':=f^{-1}(Z)$, $F:=\psi(Z)$, $Z'':=\psi^{-1}(F)$ and $F':= \rho(Z')$. Then the map $Z'\to V$ is finite and $(U,Z')$ is a Nisnevich neighbourhood of $(\bb A^1_V,F')$. Therefore the left vertical morphism in the commutative diagram
$$\xymatrix{
 H^{q}_{Z'}(U,\cal T_r(n)) \ar[r]^{}  & H^{q}_{Z''}(U,\cal T_r(n))  \\
 H^{q}_{F'}(\bb A^1_V,\cal T_r(n)) \ar[r] \ar[u]_\cong  & H^{q}_{\bb A^1_F}(\bb A^1_V,\cal T_r(n)) \ar[u]  
}$$
is an isomorphism by excision and the lower horizontal morphism is zero by Proposition \ref{proposition_A1_effacement}. This implies that the composition 
$$H^*_{Z}(W,\cal T_r(n))\to H^*_{Z'}(U,\cal T_r(n))\to H^*_{Z''}(U,\cal T_r(n))$$
is zero.

We therefore get that $(*)\colon\cal H^q_{Z^{p+1}}(\cal T_r(n))\to \cal H^q_{Z^p}(\cal T_r(n))$ is zero for all $p>0$ and it remains to show that $$(*)\colon\cal H^q_{Z^{1}}(\cal T_r(n))\to \cal H^q_{Z^0}(\cal T_r(n))$$ is zero. This map fits into the exact sequence $\cal H^q_{Z^{1}}(\cal T_r(n))\to \cal H^q_{Z^0}(\cal T_r(n))\to \cal H^q_{Z^0/Z^1}(\cal T_r(n))$. By definition $\cal H^q_{Z^0}(\cal T_r(n))_x\cong H^q_{}(\roi_{X,x}^h,\cal T_r(n))$ and $\cal H^q_{Z^0/Z^1}(n)_x\cong H^q_{}(K(\roi^h_{X,x}),\cal T_r(n))$ and we may assume that $x\in X_k$. We can therefore also show that the composition  $ H^q_{}(\roi_{X,x}^h,\cal T_r(n))\to H^q_{}((\roi_{X,x}^h)_{\eta},\cal T_r(n)) \to H^q_{}(K(\roi_{X,x}^h),\cal T_r(n))$ is injective, where $\eta$ is the generic point of the special fiber of $\Spec(\roi_{X,x}^h)$. The map $ H^q_{}(\roi_{X,x}^h,\cal T_r(n))\to H^q_{}((\roi_{X,x}^h)_{\eta},\cal T_r(n))$ is obtained by localising the exact sequence $\cal H^q_{Z^1-\eta}(\cal T_r(n))\to \cal H^q_{Z^0}(\cal T_r(n)) \to \cal H^q_{Z^0/(Z^1-\eta)}(\cal T_r(n))$ at $x$ and by what we showed above, the first map in this sequence of sheaves is zero.  
The map $H^q_{}((\roi_{X,x}^h)_{\eta},\cal T_r(n)) \to H^q_{}(K(\roi_{X,x}^h),\cal T_r(n))$ is injective by the following proposition. 

%The reduction of the case in which the residue field of $S$ is finite to the above case can be achieved by a standard norm argument. For this we refer to \cite[Proof of Thm. 2.2.7]{CHK97} and \cite[Proof of Prop. 1.9]{LuMo20}.

Finally, it follows from a norm argument
%, along the lines of the proof of the isomorphism $G(A)\to \hat{G}(A)$ for $G\in \mathfrak{CT}$ in the proof of \cite[Prop. 9]{Ke10}, 
that we can drop the assumption that the residue field of $A$ is infinite. 
We give the argument for this reduction for the injectivity following \cite[Lem. 1.9]{LuMo20}; the exactness at the other places can be shown similarly.
We only need to worry about the case that $A$ has finite residue field $k$. Let $R:=\roi_{X,x}^h$ for $x$ a point of the special fiber and let $L$ denote its residue field. Since $k$ is perfect, $L$ is a finitely generated, separable field extension of $k$ and we may therefore realise $L$ as a finite separable extension of a rational function field $k(\ul t):=k(t_1,\dots,t_d)$.
We fix a prime $p$ which is coprime to $|L:k(\ul t)|$, and let $k^{(r)}$ be the unique degree $p^r$ extension of the finite field $k$. Note that $L\otimes_kk^{(r)}=L\otimes_{k(\ul t)} k^{(r)}(\ul t)$ is the tensor product of finite field extensions of coprime degree, hence is a field.
Let $A^{(r)}$ be the finite \'etale extension of $A$ corresponding to the extension $k^{(r)}$ of $k$, and set $R^{(r)}:=R\otimes_{A}A^{(r)}$. Note that $R^{(r)}$ is local since $R^{(r)}/\mathfrak m_RR^{(r)}=L\otimes_kk^{(r)}$ is a field, and henselian since it is finite over $R$. 
%By the same argument, we may additionally assume that, denoting $R_\eta:= \roi_{X,\eta}$, the ring $R_\eta^{(r)}:= R_\eta\otimes_{A}A^{(r)}$ is local and essentially smooth.  
Consider the commutative diagram
\[
  \xymatrix{
  R^q\epsilon_*\cal T_r(n)(R^{(r)}) \ar[r]^-{} \ar@{-->}@/^5mm/[d]^N  &  R^q\epsilon_*\cal T_r(n)(\Frac(R^{(r)}))   \ar@{-->}@/^5mm/[d]^N 
  \\
R^q\epsilon_*\cal T_r(n)(R) \ar[u]  \ar[r] & \ar[u]_{} R^q\epsilon_*\cal T_r(n)(\Frac(R))  
  }
\]
in which the dotted arrows correspond to the norm map $N$ (see for example \cite[Lem. 3.4]{LuMo20} or construct the norm map using the distinguished triangle in Section \ref{section_etaletatetwists} defining $p$-adic \'etale Tate twists). Let $a$ be in the kernel of the lower horizontal map. For primes $p$ and $p'$ as above with $(p,p')=1$, by a filtered colimit argument we get that for $r$ large enough $p^ra=0=p'^ra$ and therefore $a=0$.
\end{proof}

\begin{proposition}\label{prop_effaceability} 
Let $X$ and $S$ be as above and let $x\in X$ a point of the special fiber. Then the canonical map
$$H^{q}(\roi^h_{X,x},\cal T_r(n)) \to   H^{q}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))$$
is injective for all $q\geq 0$ and $n\geq 0$.
\end{proposition}
\begin{proof}
%Let $K=\Frac(\roi^h_{X,x})$.
Let $Y_x=\Spec((\roi^h_{X,x})/\pi)$ and $i:Y_x\to \Spec(\roi^h_{X,x})$ be the inclusion. Then $c=\mathrm{codim}_{\Spec(\roi^h_{X,x})}(Y_x)=1$ and we consider the long exact localisation sequence
$$\xymatrix{ 
\dots \quad \ar[r]& H^{n-1}_{Y_x}(\roi^h_{X,x},\cal T_r(n))  \ar[r] & H^{n-1}(\roi^h_{X,x},\cal T_r(n)) \ar[r] &  H^{n-1}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))\ar[r] &  & \\
& H^{n}_{Y_x}(\roi^h_{X,x},\cal T_r(n)) \ar[r] &  H^{n}(\roi^h_{X,x},\cal T_r(n)) \ar[r] &  H^{n}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))\ar[r] & & \\
& H^{n+1}_{Y_x}(\roi^h_{X,x},\cal T_r(n)) \ar[r] &  H^{n+1}(\roi^h_{X,x},\cal T_r(n)) \ar[r] &  H^{n+1}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))\ar[r] & \quad \dots &
}$$

First note that $H^{m}_{Y_x}(\roi^h_{X,x},\cal T_r(n))=0$ for $m\leq n$ since by definition the natural map $H^{m}_{}(\roi^h_{X,x},\cal T_r(n))\to H^{m}_{}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))$ is an isomorphism for $m<n$ and $H^{n}(\roi^h_{X,x},\cal T_r(n)) \to   H^{n}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))$ is injective. By Proposition \ref{proposition_purity_tate_twists}, we have that 
$H^{n+1}_{Y_x}(\roi^h_{X,x},\cal T_r(n))\cong W_r\Omega^{n-1}_{Y_x,\log}$ and by \cite[Thm. 1.7]{LuMo20} we have an isomorphism of short exact sequences
$$\xymatrix{
0\ar[r] & \hat{K}^{n}(\roi^h_{X,x})/p^r \ar[r] \ar[d] & K^{n}(\roi^h_{X,x}[\frac{1}{\pi}])/p^r \ar[r] \ar[d]  & W_r\Omega^{n-1}_{Y_x,\log} \ar[r] \ar[d]  &0 \\
0\ar[r] & H^{n}(\roi^h_{X,x},\cal T_r(n)) \ar[r] & H^{n}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n)) \ar[r] & W_r\Omega^{n-1}_{Y_x,\log} \ar[r] &0 
}$$
implying that $H^{n+1}(\roi^h_{X,x},\cal T_r(n)) \to  H^{n+1}(\roi^h_{X,x}[\frac{1}{\pi}],\cal T_r(n))$ is injective. Finally, $$H^{m}(\roi^h_{X,x},\cal T_r(n))=0$$ for $m>n+1$. Indeed, see also the next section, by Gabber's affine analogue of proper base change \cite[Thm. 1]{GabberAffineAnalogue}, there is an isomorphism $H^{m}(\roi^h_{X,x},\cal T_r(n))\cong H^{m}(Y_x,i^*\cal T_r(n))$. But $H^{m}(Y_x,i^*\cal T_r(n))=0$ for $m>n+1$ since $\cal T_r(n)$ is concentrated in $[0,n]$ and the $p$-cohomological dimension of an affine noetherian scheme of characteristic $p$ is $\leq 1$ (see \cite[Exp. X, Thm. 5.1]{SGA4}).
\end{proof}

\begin{remark}
The relationship with \cite{SS19} is the following: Schmidt-Strunk in \textit{loc. cit.} use strict $\bb A^1$-invariance in order to obtain Lemma \ref{prop_commutativity} and therefore their results just work for such theories. Sato's $p$-adic \'etale Tate twists are not strict $\bb A^1$-invariant and we therefore use that they have a projective bundle formula in order to prove the required commutativity.
\end{remark}

\begin{remark}\label{remark_Elmanto_Morrow}
In the context of their motivic cohomology of equicharacteristic schemes, Elmanto-Morrow have introduced the term \textit{deflatable}. We first recall the definition, replacing the base field by an arbitrary base scheme $S$.
Let $\cal F:{\rm Sm}^{\rm op}_S\to \rm Sp$ be a presheaf. For $X\in {\rm Sm}_S$ we write $\cal F^X$ for the presheaf $U\mapsto \cal F(U\times_SX)$. There are morphisms of presheaves $$j^*,\pi^*\infty^*:\cal F^{\bb P^1}\to \cal F^{\bb A^1},$$
induced by the following morphisms of schemes: 
\begin{enumerate}
\item[1.] $\pi$ is the projection $\bb A^1_S\to S$,
\item[2.] $\infty$ is the closed immersion $S\to \bb P^1_S$ at infinity,
\item[3.] $j: \bb A^1_S\to\bb P^1_S$ is the open immersion complementary to the immersion in 2.
\end{enumerate}
\begin{definition}(\cite[Def. 6.7]{ElmantoMorrow})
A presheaf $\cal F:{\rm Sm}^{\rm op}_S\to \rm Sp$ is \textit{deflatable} if the maps $j^*$ and $\pi^*\infty^*$ are homotopic.
\end{definition}
Assuming that $S$ is the spectrum of a henselian dvr with perfect residue field, our arguments show 
%that any Nisnevich presheaf which satisfies the $\bb P^1$-bundle formula is deflatable and 
that any Nisnevich presheaf $\cal F:{\rm Sm}^{\rm op}_S\to \rm Sp$ which satisfies the $\bb P^1$-bundle formula is deflatable. Furthermore, if 
\begin{enumerate}
\item[1.] $\cal F$ is deflatable,
\item[2.] the natural map $\pi_q\cal F(-)\to \pi_q\cal F(-[\frac{1}{\pi}])$ is injective for all $q$,
\item[3.] $\cal F$ is finitary (i.e. commute with filtered colimits),
\item[4.] there exists a functorial norm map $N_{B'/B}:\cal F(B')\to \cal F(B)$ for any finite \'etale extension of local rings $B\to B'$ such that if $B\to B'$ is of constant degree $d$, then the composition $\cal F(B)\to\cal F(B')\xto{N_{B'/B}} \cal F(B')$ is multiplication by $d$,
\end{enumerate}
then for all integers $q$, $\pi_q\cal F(-)$ satisfies the Gersten injectivity for smooth $S$-schemes
(see \cite[Lem. 6.11]{ElmantoMorrow} in equal characteristic) and has a Gersten resolution (for more details on the Gersten resolution for spectra see \cite[Sec. 5.2]{CHK97} and \cite{SS19}) in the Nisnevich topology.
\end{remark}

\section{The $p$-adic cycle class map} 
We fix the following notation. Let $A$ be a discrete valuation ring with local parameter $\pi$, perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero. Let $S=\Spec(A)$. Let $X=\Spec(B)$ be an ind-smooth $S$-scheme. Let $I=(\pi)$ be the ideal defining the special fiber or let $I=\mathfrak{m}_B$ be the maximal ideal of $B$. We denote by $B^h$ the henselisation of $B$ along $I$. Let $i:\Spec B/I\to \Spec B^h$ be the inclusion. In the entire section let $n\geq 0$.
%Let $(B,\mathfrak{m}_B)$ be a henselian local ring of mixed characteristic $(0,p)$ which is essentially smooth over a discrete valuation ring $(A,\mathfrak{m}_A=(\pi))$ and $(p)\subset I\subset \mathfrak{m}_B$ an ideal. 
 
By Gabber's affine analogue of proper base change \cite[Thm. 1]{GabberAffineAnalogue},\footnote{If $I=\mathfrak{m}_B$ then this is easier since then the map $B^h\to B/I$ induces an equivalence between the categories of finite \'etale $B^h$-algebras and finite \'etale $B/I$-algebras and any \'etale cover of $B^h$ has a refinement by a finite \'etale cover. The statement then follows from \cite[III.3.3]{Milne1986}.} the restriction map
$$H^{s}(\Spec B^h,\T_r(n))\xto\cong H^{s}(\Spec B/I,i^*\T_r(n))$$
is an isomorphism for all $s$.
For the following theorem see also \cite[Sec. 5]{AMMN} and \cite[Thm. 3.7]{sakagaito2023etale}. In the proof we closely follow the proof given in \cite[Thm. 3.7]{sakagaito2023etale}.

\begin{theorem}\label{theorem_top_cohomology}
Let the notation be as above. Then $R\epsilon_* i^*\T_r(n)\in D((\Spec B/I)_{\rm Nis},\bb Z/p^r)$ is concentrated in degree $[0,n+1]$. Furthermore,
$$H^{n+1}(\Spec B^h,\T_r(n))\cong H^{1}(\Spec B/I,W_r\Omega^n_{\log}).$$
\end{theorem}
\begin{proof}
The first statement follows from 
%the spectral sequence $$E_2^{s,t}=H^s_{\rm Nis}(\Spec B/I,R^t\epsilon_*(i^*\T_r(n)))\Longrightarrow H^s_\et(\Spec B/I,i^*\T_r(n)),$$ 
Property (ii) of $\T_r(n)$ and the fact that the $p$-cohomological dimension of an affine noetherian scheme of characteristic $p$ is $\leq 1$ (see \cite[Exp. X, Thm. 5.1]{SGA4}).
We turn to the second statement. By the above, it suffices to show that $$H^{n+1}(\Spec B/I,i^*\T_r(n))\cong H^{1}(\Spec B/I,W_r\Omega^n_{\log}).$$
The commutative diagram of distinguished triangles
$$\xymatrix{
i^*\T_r(n)_X \ar[r]^{} \ar[d]_{} & i^*\tau_{\leq n}Rj_*\mu_{p^r}^{\otimes n}\ar[d]^{}  \ar[r]  & W_r\Omega^{n-1}_{B/I\log}[-n] \ar@{=}[d]_{} \\
FM^n_r \ar[r]  &  i^*R^nj_*\mu_{p^r}^{\otimes n}[-n] \ar[r]^{} & W_r\Omega^{n-1}_{B/I\log}[-n], 
}$$
in which $FM^n_r$ is defined to be the kernel of the lower right horizontal map, gives an isomorphism
$$H^{n+1}(\Spec B/I,i^*\T_r(n))\cong H^{1}(\Spec B/I,FM^n_r).$$
The unit filtration gives an exact sequence 
$$0\to U^1M^n_r\to FM^n_r\to W_r\Omega^n_{\log}\to 0$$
(see \cite[Thm. 3.4.2]{Sa07}).
%We assume now for simplicity that $p>2$. We first show the case $r=1$, the case $r>1$ follows from a Bockstein sequence argument. 
If $r=1$, then $U^1M^n_r$ is a coherent module given in terms of differential forms (see \cite[Cor. 1.4.1]{BlochKato1986}) and therefore its higher cohomology groups vanish on the affine scheme $\Spec B$. This implies that 
$$H^{n+1}(\Spec B/I,i^*\T_1(n))\cong H^{1}(\Spec B/I,FM^n_1)\cong H^{1}(\Spec B/I,W_1\Omega^n_{\log}).$$
The case $r>1$ follows by induction and the five-lemma from the commutative diagram
$$\xymatrix{
  & H^{n+1}(\Spec B/I,i^*\T_r(n)) \ar[r]^{} \ar[d]_{\cong} & H^{n+1}(\Spec B/I,i^*\T_{r+1}(n)) \ar[d]^{}  \ar[r]  & H^{n+1}(\Spec B/I,i^*\T_1(n)) \ar[d]_{\cong} \ar[r]  & 0 \\
0\ar[r]  & H^{1}(\Spec B/I,W_r\Omega^n_{\log}) \ar[r]  &  H^{1}(\Spec B/I,W_{r+1}\Omega^n_{\log}) \ar[r]^{} & H^{1}(\Spec B/I,W_1\Omega^n_{\log}). & 
}$$

The upper sequence is exact by \cite[Prop. 4.3.1]{Sa07} and the vanishing of $H^{n+2}(\Spec B/I,i^*\T_1(n))$. The lower sequence is exact by \cite[p. 779, Lem. 3]{CSS83} and the the surjectivity of $W_{r+1}\Omega^n_{\log}(\Spec B/I)\cong\cal K^M_{n}(\Spec B/I)/p^{r+1}\to K^M_{n}(\Spec B/I)/p\cong W_1\Omega^n_{\log}$.
\end{proof}
%\section{Kato conjecture over a DVR with $p$-coefficients}
%Use rigidity of $\cal H^{n+1}(\T(n))$.

\begin{theorem}\label{theorem_comparison}
%Let $A$ be a discrete valuation ring with local parameter $\pi$, perfect residue field $k$ of characteristic $p>0$ and fraction field $K$ of characteristic zero. 
Assume that $k$ is finite. Let $X$ be a smooth projective $S$-scheme with special fiber $X_k$ (in the following also denoted by $X_1$). We denote the inclusion of the special fiber by $i:X_k\to X$ and set $X_n=X\times_A A/\pi^n$. Let $d$ be the relative dimension of $X$ over $S$.
Then for $q\in\{0,1\}$ there are isomorphisms
$$H^{d-q}_{\rm Nis}(X_k,i^*\hat{\cal K}^M_{d,X}/p^r) \xto{\cong}  \projlim_n H^{d-q}_{\rm Nis}(X_k,\hat{\cal K}^M_{d,X_n}/p^r)\xto{\cong} H^{2d-q}_\et(X_k,i^*\cal T_r(d)). $$
\end{theorem}
\begin{proof}
%The Galois symbol map $\hat{\cal K}^M_{d,X}/p^r\to \cal H^d(\cal T_r(d)_X)$ induces a map $cl_X:H^d(X,\hat{\cal K}^M_{d,X}/p^r)\to H^{2d}(X,\cal T_r(d))$. 
%Let $R\epsilon_*\cal T_r(d)_{X}:=\cal T_r(d)_{X,\rm Nis}$.
In the following we denote the change of sites $X_\et\to X_{\rm Nis}$ as well as the change of sites $({X_k})_\et\to ({X_k})_{\rm Nis}$ by $\epsilon$. We first note that the Galois symbol map $i^*\hat{\cal K}^M_{d,X,\rm Nis}/p^r\to i^*\cal H^d(\cal T_r(d)_{X,\rm Nis})$ factorises through $\K^M_{d,X_n,\rm Nis}/p^r$ for $n$ large enough. Via a norm argument one first reduces this statement to the case in which the residue field is large and therefore the improved Milnor $K$-theory coincides with ordinary Milnor $K$-theory. This case can be deduced from the short exact sequence $0\to 1+\pi^n\roi_X \to \roi_X^\times \to \roi_{X_n}^\times \to 0$, which induces the unit filtration on Milnor $K$-theory, noticing that in the Nisnevich topology the group $(1+\pi^n\roi_X)/p^r=0$ for $n$ large enough by Hensel's lemma \cite[II, Lem. 2]{Elkik1973}. 
Indeed, the kernel of the restriction map $i^*\cal K^M_{d,X}\to \cal K^M_{d,X_n}$ is locally generated by elements of the form $\{1+\pi^n a,f_1,\dots,f_{d-1}\}$ with $f_1,\dots,f_{d-1}$ units \cite[Lemma 1.3.1]{KaS86}. This implies the first isomorphism of the theorem (without any restriction on $q$).
Next we consider the Leray spectral sequence
$$H^p_{\rm Nis}(X_k,R^q\epsilon_*(i^*\cal T_r(d)_X))\Longrightarrow H^{p+q}_\et(X_k,i^*\cal T_r(d)).$$
By \cite[Prop. 2.2(b)]{Geisser2004} we have that $R^q\epsilon_*(i^*\cal T_r(d)_X)\cong i^*R^q\epsilon_*(\cal T_r(d)_{X})$.
By \cite{LuMo20} the map $i^*\hat{\cal K}^M_{d,X,\rm Nis}/p^r\to  i^*R^d\epsilon_*(\cal T_r(d)_{X})$ is an isomorphism and by Theorem \ref{theorem_top_cohomology} we know that $R^{d+1}\epsilon_*(i^*\cal T_r(d)_X)= R^1\epsilon_*W_r\Omega^d_{\log}$ and $R^q\epsilon_*(i^*\cal T_r(d)_X)=0$ for $q>d+1$. 
This implies the existence of an exact sequence %\morten{Ist wirklich $H^{d-1}_{\rm Nis}(X_k,i^*\hat{\cal K}^M_{d,X}/p^r)\to H^{2d-1}_{\rm Nis}(X_k,i^*\cal T_r(d))$?}
$$H^{d-3}_{\rm Nis}(X_k,R^1\epsilon_*W_r\Omega^d_{\log})\to H^{d-1}_{\rm Nis}(X_k,i^*\hat{\cal K}^M_{d,X}/p^r)\to H^{2d-1}_\et(X_k,i^*\cal T_r(d))\to $$
$$H^{d-2}_{\rm Nis}(X_k,R^1\epsilon_*W_r\Omega^d_{\log})\to H^{d}_{\rm Nis}(X_k,i^*\hat{\cal K}^M_{d,X}/p^r)\to H^{2d}_\et(X_k,i^*\cal T_r(d))\to $$
$$H^{d-1}_{\rm Nis}(X_k,R^1\epsilon_*W_r\Omega^d_{\log})\to 0\to H^{2d+1}_\et(X_k,i^*\cal T_r(d)) \to  $$
$$H^{d}_{\rm Nis}(X_k,R^1\epsilon_*W_r\Omega^d_{\log})\to 0.$$
For the exactness of the first row note that the Gersten resolution implies that $H^{d}_{\rm Nis}(X_k,R^{d-1}\epsilon_*(i^*\cal T_r(d)_X))=0$. The Kato conjectures say that $H^{q}_{\rm Nis}(X_k,R^1\epsilon_*W_r\Omega^d_{\log})=0$ for $q<d$ and isomorphic to $\bb Z/p^r$ for $q=d$. By \cite[Thm. 0.3]{JS} this holds for $d-4\leq q \leq d$ which implies the theorem.
\end{proof}

Let $q\in\{0,1\}$. Theorem \ref{theorem_comparison} implies that there is a commutative diagram
$$
\xymatrix{
H^{d-q}_{\rm Nis}(X,\hat{\cal K}^M_{d,X}/p^r)\ar[r]^-{res^d_{/p^r}} \ar[d]_{cl_X} & \projlim_n H^{d-q}_{\rm Nis}(X_k,\K^M_{d,X_n}/p^r) \ar[d]^{cl_{X_k}}  \\
H^{2d-q}_\et(X, \cal T_r(d)) \ar[r]^-\cong & H^{2d-q}_\et(X_k,i^*\cal T_r(d))   \\
}
$$
in which the bottom map is an isomorphism by \'etale proper base change and $cl_{X_k}$ is an isomorphism by Theorem \ref{theorem_comparison}. 
By \cite[Thm. 0.2]{LuMo20} there is an isomorphism 
$\CH^d(X)/p^r\cong H^{d}_{\rm Nis}(X,\hat{\cal K}^M_{d,X}/p^r)$.
By \cite[Proof of Cor. 2.7]{Lu16} and the Gersten conjecture for $\hat{\cal K}^M_{d,X}/p^r$ \cite[Thm. 0.2]{LuMo20}, there is an isomorphism
$\CH^d(X,1,\bb Z/p^r)\cong H^{d-1}_{\rm Nis}(X,\hat{\cal K}^M_{d,X}/p^r).$
It is an interesting question to ask if the maps $res^d_{/p^r}$ and $cl_X$ are isomorphisms (see \cite{Lueders2019}, \cite{SS14} for the surjectivity and \cite[Conj. 10.1]{KerzEsnaultWittenberg2016} for the original question in the case $q=0$).

\bibliographystyle{acm}
\bibliography{Bibliografie} 

\noindent
\parbox{0.5\linewidth}{
\noindent
Morten L\"uders \\
Universität Heidelberg\\
Mathematisches Institut \\
Im Neuenheimer Feld 205 \\
69120 Heidelberg \\
Germany\\
{\tt mlueders@mathi.uni-heidelberg.de}
}
\end{document}